OpenStudy (jollypikachu300):

WILL FAN AND MEDAL I NEED HELP There are 150 trees in the park. Mark noticed that 120 of them are maple trees. What percent of the trees are maple? A. 30% B. 80% C. 120% D. 150%

OpenStudy (anonymous):

It is A

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

eliminate c and d and a is too small so xD

OpenStudy (anonymous):

Sorrty he has listed A and B as 80%

OpenStudy (liamschumm):

So, here is a great formula:

OpenStudy (liamschumm):

Number/Total

OpenStudy (liamschumm):

That gives what percent number is of total!

OpenStudy (jollypikachu300):

wouldn't the answer be 30% because I subtracted

OpenStudy (liamschumm):

Whoa, whoa!

OpenStudy (liamschumm):

Hold on!

OpenStudy (liamschumm):

Number/Total.

OpenStudy (liamschumm):

120/150...

OpenStudy (anonymous):

It's 80%. 120/150 = 0.8 0.8*100 = 80

OpenStudy (liamschumm):

Cancel the 0s...

OpenStudy (jollypikachu300):

ok

OpenStudy (liamschumm):

12/15... 3*4/3*5... 4/5 = 0.8. So 80%.

OpenStudy (liamschumm):

Also, does 30% sound right?